OpenStudy (anonymous):

what are the zeros of the graph y=2x^2 -x -15?

OpenStudy (johnweldon1993):

Have you heard of the quadratic formula?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

So you have an equation 2x² - x - 15 let your coefficient correspond to the different letters a = your first coefficient = 2 b = your second coefficient = -1 c = your third coefficient = -15 plug them into that formula \[\frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] This will tell you the zeros (x-intercepts) of the graph

OpenStudy (anonymous):

sense its already -1 would it change to poitive1?

OpenStudy (johnweldon1993):

I assume you're talking about the -b part? then yes you would

OpenStudy (anonymous):

ya so the answer would be 2.81 & -2.56?

OpenStudy (johnweldon1993):

\[\frac{ 1 \pm \sqrt{-1^2 - 4(2)(-15)} }{ 2(2) }\] \[\frac{ 1 \pm \sqrt{1 + 120 } }{ 4 }\] \[\frac{ 1 \pm \sqrt{121} }{ 4 }\] \[\frac{ 1 \pm 11 }{ 4 }\] so breaking that into the 2 equations \[\frac{ 1 + 11 }{ 4 } =?\] and \[\frac{ 1 - 11 }{ 4 } = ?\]

OpenStudy (anonymous):

how did you get 121? I redid it and got 119

OpenStudy (johnweldon1993):

remember...-1 TIMES -1 = 1 right? you have -1 times -1 = 1 -4 times 2 times -15 = 120 1 + 120 = 121

OpenStudy (anonymous):

so its not just -1^2-4(2)(-15)?

OpenStudy (johnweldon1993):

late response \[\sqrt{-1^2 - 4(2)(-15)}\] \[\sqrt{1 - (-120)}\] \[\sqrt{121} = 11\]

OpenStudy (anonymous):

I didn't learn it that way

OpenStudy (johnweldon1993):

Well that was just 1 portion of the quadratic equation \[\frac{ 1 \pm \sqrt{-1^2 - 4(2)(-15)} }{ 2(2) }\] how did you learn it?